# Bayesian dynamic borrowing considering semantic similarity between outcomes for disproportionality analysis in FAERS

### v2

@@ Abstract
We present a Bayesian dynamic borrowing (BDB) approach to enhance the quantitative identification of adverse events (AEs) in spontaneous reporting systems (SRSs). The method embeds a robust meta-analytic predictive (MAP) prior within a Bayesian hierarchical model and incorporates semantic similarity measures (SSMs) to enable weighted information sharing from MedDRA Preferred Terms (PTs) that are clinically similar to the target PT. This continuous similarity-based borrowing addresses limitations of rigid hierarchical grouping in current disproportionality analysis (DPA).
Using data from the FDA Adverse Event Reporting System (FAERS) between 2015 and 2019, we evaluate this approach - termed IC SSM - against standard Information Component (IC) analysis and IC with borrowing at the MedDRA high-level group term (HLGT) level. A novel references set (PVLens), derived from FDA product label updates, enabled prospective evaluation of method performance in identifying AEs prior to official labeling.
The IC SSM approach demonstrated improved sensitivity compared to both traditional IC and HLGT-based borrowing, with minor trade-offs in F1 scores and Youden’s index. IC SSM consistently identified more true positives and detected signals over 5 months sooner than traditional IC. Despite a marginally lower aggregate Youden’s index, IC SSM showed higher performance in the early post-marketing period, providing more stable and relevant estimates than HLGT-based borrowing and traditional IC.
These findings support the use of SSM-informed Bayesian borrowing as a scalable and context-aware enhancement to traditional DPA methods. Future research should validate this approach across other datasets and explore additional similarity metrics and Bayesian inference strategies using case-level data.
@@ 1 Introduction
Quantitative analysis of ICSRs was first proposed in the late 1960s and 1970s [3, 4]. Although early applications were sporadic and often limited to specific product issues [5, 6], the 1990s marked a turning point, with routine use of these methods driven by increasing data availability and advances in computation [7]. While some quantitative signal detection approaches incorporate external data such as sales or prescription volumes to construct background rates [8, 9], the most widely used methods rely solely on ICSR data due to their practicality and timeliness [10, 11, 12, 13, 14, 15, 16] [^1] It should be noted that sales and prescription data are routinely used in other contexts, such as to establish denominators in PSURs and PBRERs; the distinction here is that they are less commonly used directly in disproportionality-based signal detection.. The standard methods—introduced in the late 1990s and early 2000s—are based on two-by-two contingency tables and remain the foundation of quantitative signal detection [17, 18, 19, 20], despite their simplicity [21]. Although more complex techniques have since been proposed, they have not consistently outperformed these foundational methods in real-world settings [22]. Notably, a multi-database comparison from the IMI PROTECT project showed that variation in signal detection performance across methods was often more influenced by the chosen detection threshold than by the specific algorithm employed [23].
@@ 2.3 Statistical methods
$y$ not $y$
$x$ $a$ $b$
not $x$ $c$ $d$
Table 1: Contingency table for product–event pairs
@@ 3.2 Performance analyses over the entire study period
IC IC HLGT
IC SSM $+$ -- $+$ --
$+$ 1,167 165 1,259 73
-- 3 1,002 41 964
Table 2: Two-by-two contingency tables comparing IC SSM with IC and IC HLGT for detecting positive controls
@@ 5 Conclusion
This study highlights the potential of semantic similarity-informed borrowing to improve post-marketing safety surveillance through more nuanced DPA. Compared to the standard IC approach – routinely used in PV – and borrowing at the MedDRA HLGT level, the IC SSM method has the ability to detect true positive signals earlier, with improved sensitivity and favorable trade-offs in F1 score and Youden’s index.
@@
6 Declarations
GSK covered all costs associated with the conduct of the study and the development of the manuscript and the decision to publish the manuscript. J.P., F.H., G.P, A.C., and A.B. are employed by GSK and hold financial equities. This manuscript has not been submitted to, nor is under review at, another journal or other publishing venue. The authors have no competing interests to declare that are relevant to the content of this article.
Data availability All data and code to support this analysis is available from https://github.com/jlpainter/BDB.
Supplement

### v3

# Semantic Similarity-Informed Bayesian Borrowing for Quantitative Signal Detection of Adverse Events

@@ Abstract
We present a Bayesian dynamic borrowing (BDB) approach to enhance the quantitative identification of adverse events (AEs) in spontaneous reporting systems (SRSs). The method embeds a robust meta-analytic predictive (MAP) prior within a Bayesian hierarchical model and incorporates semantic similarity measures (SSMs) to enable weighted information sharing from clinically similar MedDRA Preferred Terms (PTs) to the target PT. This continuous similarity-based borrowing overcomes limitations of rigid hierarchical grouping in current disproportionality analysis (DPA).
Using data from the FDA Adverse Event Reporting System (FAERS) between 2015 and 2019, we evaluate our approach - termed IC SSM - against traditional Information Component (IC) analysis and IC with borrowing at the MedDRA high-level group term level (IC HLGT). A reference set (PVLens), derived from FDA product label updates, enabled prospective evaluation of method performance in identifying AEs prior to official labeling.
The IC SSM approach demonstrated higher sensitivity (1332/2337=0.570, Youden’s J=0.246) than traditional IC (Se=0.501, J=0.250) and IC HLGT (Se=0.556, J=0.225), consistently identifying more true positives and doing so on average 5 months sooner than traditional IC. Despite a marginally lower aggregate F1-score and Youden’s index, IC SSM showed higher performance in early post-marketing periods or when the detection threshold was raised, providing more stable and relevant alerts than IC HLGT and traditional IC.
These findings support the use of SSM-informed Bayesian borrowing as a scalable and context-aware enhancement to traditional DPA methods, with potential for validation across other datasets and exploration of additional similarity metrics and Bayesian strategies using case-level data.
@@ 1 Introduction
Quantitative analysis of ICSRs was first proposed in the late 1960s and 1970s [3, 4]. Although early applications were sporadic and often limited to specific product issues [5, 6], the late 1990s marked a turning point, with routine use of these methods driven by increasing data availability and advances in computation [7]. While some quantitative signal detection approaches incorporate external data such as sales or prescription volumes to construct background rates [8, 9], the most widely used methods rely solely on ICSR data due to their practicality and timeliness [10, 11, 12, 13, 14, 15, 16] [^1] It should be noted that sales and prescription data are routinely used in other contexts, such as to establish denominators in PSURs and PBRERs; the distinction here is that they are less commonly used directly in disproportionality-based signal detection.. The standard methods—introduced in the late 1990s and early 2000s—are based on two-by-two contingency tables and remain the foundation of quantitative signal detection [17, 18, 19, 20], despite their simplicity [21]. Although more complex techniques have since been proposed, they have not consistently outperformed these foundational methods in real-world settings [22]. Notably, a multi-database comparison from the IMI PROTECT project showed that variation in signal detection performance across methods was often more influenced by the chosen detection threshold than by the specific algorithm employed [23].
@@ 5 Conclusion
This study shows that semantic similarity-informed borrowing can improve post-marketing safety surveillance by enabling more nuanced disproportionality analysis (DPA). Compared to the standard IC approach – routinely used in PV – and borrowing at the MedDRA HLGT level, the IC SSM method has the ability to detect true positive signals earlier, with improved sensitivity and favorable trade-offs in F1 score and Youden’s index.
@@
Declarations
Author contributions
François Haguinet and Jeffery Painter contributed to the conceptual design of the study, authored the manuscript, and participated in its critical revision. Gregory Powell provided clinical expertise and reviewed and edited the manuscript. Andrea Callegaro contributed to the statistical methodology and reviewed and edited the manuscript. Andrew Bate brought forward the idea of the new approach, contributed to the conceptual design, and reviewed and edited the manuscript.
Financial disclosure
GSK covered all costs associated with the conduct of the study and the development of the manuscript and the decision to publish the manuscript. J.P., F.H., G.P, A.C., and A.B. are employed by GSK and hold financial equities.
Appendix